OpenStudy (esraa):

When a capacitor has a charge of absolute value 80 µC on each plate the potential difference across the plates is 10 V. How much energy is stored in this capacitor when the potential difference across its plates is 120 V?

OpenStudy (unklerhaukus):

The definition of capacitance C is the amount of charge Q that can be stored for a given potential V_ab.\[C=\frac QV_{ab}\]

OpenStudy (unklerhaukus):

A certain capacitor will have a constant capacitance

OpenStudy (anonymous):

So energy=1/2CV^2 or 1/2QV so CV=Q C * 10=80*10^-6 C=8*10^-6 1/2 * 8*10^-6 * 120^2 =E .0576J

OpenStudy (marco26):

\[C = q/v = 80x10^-6/10 = 8x10^-6\]\[W= 1/2 (C)(V^2) = 1/2 (8x10^-6)(120)^2= 0.0576 J\]

OpenStudy (anonymous):

.576J=57.6mj :D

OpenStudy (esraa):

Thnx alot :D